Phone Numbers

March 25, 1988

Name  location  Number
Dr. Andy Boughton Tech Square (NE43-255)  x3-8835
Ike Chuang Conner x5-8207
Hand Lab (9th floor)  x3-2077
André DeHon East Campus x5-6369
Tech Square (NE43-226)  x3-2795
CSG Hardware Lab  x3-0794
Brian LaMacchia MacGregor x5-9210
Tech Square (NE43-428)  x3-8570
TMRC  x3-3269
Debby Wallach Conner x5-8213
Tech Square (NE43-521c) x3-6204
Stan Zanarotti Tang 494-0170
Tech Square (NE43-510)  x3-6042
Watchmaker Area  E40 3rd floor  x3-4261
Paul Willis  Home  782-9697
Brian Avery  Student House (?)  734-5974
Jon Freeman  Next House  x5-8718
Next House Desk  x3-8761
office  x3-7178
Trang Hoang  BU (dorm)  375-6518
Stanley Salters  Berklee  267-9493

98 Hemingway St.

Jenifer Tidwell Pika (institute) x3-5009
Pika (Cambridge)  492-6983
Mike Ward  Kappa Sigma  x5-9546